Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

2021-02-02

Observers:   Jacco Terwel, Mark Magee
Support: Simon Prentice, Luke Harvey
TAT: Michael Fulton, Shubham Srivastav, Janet Chen

EFOSC Calibrations: bias, calib, grz sky flats


Weather:
Start: wind 3.2, humidity 39.0, seeing
Middle: wind 2.6, humidity 27.0, seeing 2.3
End:

Comments: Seeing was very erratic during the night. Usually between 1.5 - 2. Somewhere around ~>30 minutes were lost due to a problem with the filter wheel.

Started with EFOSC
 UT Target Type Set up
 Slit Exposure Airmass Seeing Notes
00:45
 AT2020adkm Follow up
 gr131.0
2700s
1.5

Plan is for second exposure tomorrow
 01:35 GD71 Standard gr11,13,16 1.0 100s 1.4 1.5 
 01:56 SN2020vef Follow up
 gr11 1.0 2700s 1.5 1.5Seeing jumped up to 1.8 in the middle of the exposure
 02:51 SN2020znl Follow up
 gr11 1.5 2700s 1.06 1.7 
 03:50        Focus
 04:00 AT2021bmt Classification gr13 1.5 900s 1.1 1.7 Ia
 04:23 ATLAS21cbe Classification gr13 1.5 900s 1.1 2.2Observed last night but we took another spectrum, looks like just a galaxy?
 04:55 AT2021bmc Classification gr13 1.5 900s 1.4 1.5 Blue, some weak and broad features?
 05:15 L745A Standard gr11,13,16 1.5 100s 1.1 1.5 

 05:30        Problem with the filter wheel, had to be reset
 06:02 AT2021bma Classification gr13 1.5 600s 1.2 1.2 Looks like junk. Very noisy, but we think the right object was in the slit
 06:20     SN2020uik Follow up
      Couldn't see anything in 60s acquisition. Aborted
 06:25 ATLAS21dkt Classification gr13 1.5 300s 1 1.4 Ia
 06:55 SN2021zj Follow up
 gr11 1.5 1800s 1.9 1.3 Check that the right object was in the slit...
 07:40 AT2021ble Classification gr13 1.5 900s 1.1 1.3 Ia - young
 08:01        Focus
 08:09 LTT3864 Standard gr11,13,16 1.5 100s1.1
 1.3 
 08:27 AT2021bmf Classification gr13 1.5 600s 1.8 1.6 Ic-BL?
 09:00 AT2020acka Classification gr13 1.5900s
 2 1.7 
Comments